OpenStudy (anonymous):

What must be added to 3x -7 to make x2 + 4x -1?

OpenStudy (anonymous):

Okay, here subtract the second equation by the first and simplify

OpenStudy (anonymous):

Are you having problems with simplification?

OpenStudy (anonymous):

I see an answer coming out there.

Directrix (directrix):

Let y represent what must be added Then, y + 3x -7 = x^2 + 4x -1 y = x^2 + 4x -1 - 3x - (-7) y = x^2 + x + 6 Check by adding (x^2 + x + 6) to 3x - 7 to determine if the sum is x^2 + 4x -1. @RyanTLopez